OpenStudy (anonymous):

A line passes through (1, -4) and has a slope of 3. What is an equation of the line in point-slope form?

jimthompson5910 (jim_thompson5910):

y - y1 = m(x - x1) ... point-slope form y - y1 = 3(x - x1) ... plug in the given slope m = 3 y - (-4) = 3(x - 1) ... plug in the given point (1,-4) y + 4 = 3(x - 1) ... simplify